# Generalized Hypercube Queuing Models with Overlapping Service Regions

### v2

@@ 3.1 Synthetic results
In this section, we delve into the assessment of the efficacy of the proposed hyperlattice queuing models via synthetic experiments. These experiments employ diverse parameter settings within synthetic service systems to investigate the accuracy of the performance metrics estimated by our models. To embark on this, we first use the hyperlattice queuing model to estimate the steady-state probabilities of the systems by solving the balance equations (4). Based on the solutions, we then calculate summary statistics, including individual workloads and the fraction of dispatches, according to (11) and (12). To validate the accuracy of our model estimations, we also developed a simulation program using simpy [17], a process-based discrete-event simulation framework based on Python. The primary purpose of this simulation program is to provide a ground truth of the queuing dynamics of the system so that we can compare the estimated performance measures from our model with the simulation results. Each simulation generates 1,000 events, and servers are dispatched from mobile positions. If a server is idle, its position is selected from a uniform distribution over its service region. Otherwise, the server departs from the location of the previous call it is responding to. The servers’ travel speed is a constant value $c=10$. Using the simulation results, we can approximate the true individual workload of each server $\rho_{i}$ by calculating the percentage of time that the server is busy, and the true fraction of dispatches $\rho_{i,e}$ (or $\rho_{i,i}$) by calculating the percentage of calls received in region $e$ (or $i$) that have been assigned to server $i$.
@@ 3.1.1 Overlapping patrolling
We also test our models on the synthetic systems that allow overlapping patrolling. Specifically, we create a two-server system where the service region is a 100 by 100 square with a single overlapping service region in the middle ($I=2$ and $E=1$), as depicted in Figure 5 (a). We assume that calls are uniformly distributed with an arrival rate of $\lambda=1$ and that the service rate of each server is fixed at $\mu_{i}=1$. Changing the hyperparameters of the system, such as the size of the overlapping region and the dispatch policy, can significantly shift the queuing dynamics. Our numerical experiments suggest that the hyperlattice queuing models can accurately capture intricate queuing dynamics under different hyperparameter settings for such a system.
(a) $r=0$
(b) $r=0.4$
(c) $r=0.7$
(d) Individual workloads
(e) Fraction of dispatches
Figure 13: Synthetic results for two-server systems with varying overlapping ratio $r$. In (a), (b), and (c), the shaded area represents the overlapping region of the system. (d) and (e) show individual workloads $\rho_{i}$ and fraction of dispatches $\rho_{i,(i,j)}$ with varying overlapping ratio $r$. The solid lines represent the estimated performance measures using hyperlattice queuing models, whereas the dashed lines represent the true performance measures obtained from simulation runs.
To examine the impact of overlapping ratio on system dynamics, we first create 100 systems by varying the overlapping ratio $r$. This ratio represents the percentage of the area of the overlapping region with respect to the area of the entire service region of server 1, i.e., $r=|\mathcal{O}_{\{1,2\}}|/(|\mathcal{O}_{\{1,2\}}|+|\mathcal{P}_{1}|)$, as shown in Figure 13 (a-c). We then compare the estimated values of $\rho_{i}$ and $\rho_{i,e}$ to their true values (indicated by dashed lines) in Figure 13 (d) and (e), and find that the estimated values are a good approximation of their true counterparts.
(a) $\eta_{(1,2)}(1)=0$ and $\eta_{(1,2)}(2)=1$
@@
Figure 14: Synthetic results for two-server systems with varying dispatch policy $\eta_{(1,2)}(1)$ and $\eta_{(1,2)}(2)$. In (a), (b), and (c), the color depth of region $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ represents the value of $\eta_{ij}$ and $\eta_{21}$, respectively. (d) and (e) show individual workloads $\rho_{i}$ and fraction of dispatches $\rho_{i,(i,j)}$ with varying dispatch policy $\eta_{(1,2)}(1)$ and $\eta_{(1,2)}(2)$. The solid lines represent the estimated performance measures using hyperlattice queuing models, whereas the dashed lines represent the true performance measures obtained from simulation runs.
@@
Furthermore, we also investigate the impact of dispatch policy by creating another 100 systems with different values of $\eta_{(1,2)}(1)$ and $\eta_{(1,2)}(2)$ (where $\eta_{(1,2)}(1)=1-\eta_{(1,2)}(2)$), while keeping the service regions the same, as shown in Figure 14 (a-c). For the sake of simplicity, we here omit the subscript $u$ in the notation of dispatch policy $\eta_{e}(i),~{}\forall i\in e$ as we assume it doesn’t depend on the state $u$ in this experiment. Figure 14 (d) and (e) show that the estimated $\rho_{i}$ and $\rho_{i,e}$ (indicated by solid lines) are able to closely approximate their true values suggested by the simulation.
Figure 15: An illustration depicting the experimental setup for the simulation. The dots indicate potential call reception locations, while the dashed lines illustrate the servers’ travel paths during the simulations.
3.1.2 Comparison with Hypercube queue
As outlined in Section 1, a core objective in formulating the hyperlattice model is to extend the classic hypercube queue model to more accurately represent heavy-traffic situations, thus allowing for a detailed analysis of oversaturated systems with queue lengths greater than one. In pursuit of this, we construct a synthetic system subjected to a traditional non-overlapping districting, defined by a $2\times 2$ grid encompassing four identical sub-regions served by their respective servers, as shown in Figure 15. For a meaningful comparison with the hypercube queuing model – where servers can attend to calls from any region if other servers are unavailable – we introduce an overlapping service zone comprised of all four sub-regions. This design ensures no primary service area is designated, granting every server the flexibility to address calls from any sub-region. Furthermore, we operate under the assumption that the call distributions are even across these sub-regions. This results in a call arrival rate of $\lambda/4$ = 1 for each. All servers in this model are consistent in their performance, exhibiting a uniform service rate denoted as $\mu$ = 1. Lastly, we adopt a standardized dispatch policy that is uniformly applied across all servers.
(a) Individual workloads
(b) Fraction of dispatches
Figure 16: Synthetic results for four-server systems with varying $\lambda/\mu$ ratio. Figures (a) and (b) show individual workloads $\rho_{i}$ and fraction of dispatches to its own region $\rho_{i,i}$ with varying $\lambda/\mu$ ratio. The solid lines represent the estimated performance measures using hyperlattice queuing models, whereas the dashed lines represent the true performance measures obtained from simulation runs.
(a) Individual workloads
(b) Fraction of dispatches
Figure 17: Frobenius norm of the difference of individual workloads $\rho_{i}$ and fraction of dispatches $\rho_{i,e}$ derived from the hypercube/hyperlattice queuing models and simulation results (denoted as $\hat{\rho}_{i}$ and $\hat{\rho}_{i,e}$) for the systems with varying $\lambda/\mu$ ratio. The dashed lines and the color regions represent the mean and the 95% confidence interval of the norm obtained from 10 simulation runs.
Figure 17 demonstrates the discrepancies in the estimation of individual workloads and dispatch fractions between the hypercube and hyperlattice models. In Figure 17 (a), it is evident that the hypercube model’s effectiveness diminishes substantially as the ratio escalates, whereas our method sustains commendable performance even at higher ratios. This advantage is primarily attributed to the hyperlattice’s capability to account for more detailed dynamics between queues, a feature not present in the hypercube model. Figure 17 (b) further reinforces the superior predictive accuracy of the hyperlattice model in estimating the fraction of dispatches, regardless of the $\lambda/\mu$ ratio. The success is attributed to the harmonized policy adherence between the simulation and that implemented by the hyperlattice model. Furthermore, this dispatch policy demonstrates robustness against fluctuations in the $\lambda/\mu$ ratio. Conversely, the hypercube model, which assumes a shared queue among all servers, fails to accurately capture heavy traffic dynamics due to its oversimplified representation of the service process. Moreover, the hypercube model is constrained to a myopic policy (usually only dependent on travel distance) adopting a fixed preference for dispatch that becomes noticeably inadequate at lower $\lambda/\mu$ ratios, leading to a pronounced divergence from simulation results. However, an increase in the ratio sees a shift in the hypercube model towards a homogenous policy, akin to the one employed in our simulation, consequently reducing the gap in estimation accuracy. Such shift underscores the limitations in the hypercube model, demonstrating its restricted adaptability to different scenarios.
@@ 3.2 Case study: Police Redistricting in Atlanta, GA
Figure 18: Figure (a) presents a comprehensive map delineating the various police zones in Atlanta, Georgia, while Figure (b) specifically focuses on Zone 5. The intensity of color shading directly corresponds to the frequency of police call arrivals, providing a visual representation of call distribution across the zones.
@@
In this section, we present a case study focusing on the police redistricting problem within Atlanta, Georgia. Figure 18 illustrates the various police zones of the city, with a particular emphasis on the intricate beats of Zone 5. Ideally, each beat would be served by a dedicated police unit, ensuring a focused response within its boundaries. However, this is often not feasible due to a limited number of available police units, primarily attributed to constraints in officer and resource allocation. In instances where certain beats lack a dedicated police unit, units from adjacent beats are required to provide assistance. While such support is permissible among units within the same zone, assistance across zones is explicitly prohibited. The depth of color shading in Figure 18 is a direct representation of the frequency of police call arrivals, quantified by the arrival rate $\lambda$. These rates are estimated based on an analysis of historical 911 calls-for-service data, specifically collected during the years 2021 and 2022, as detailed in the study of [25]. The color reveals that Zone 5 experiences the highest frequency of police calls, approaching an average of 18 calls per hour. As a result, our analysis concentrates on examining potential overlapping districting plans specific to the beats in Zone 5. A key performance metric under consideration is the standard deviation of the individual workloads, as previously discussed in Section 2.4.1.
(a) $\sigma$ = 0.13
(b) $\sigma$ = 0.11
(c) $\sigma$ = 0.08
(d) $\sigma$ = 0.13
(e) $\sigma$ = 0.10
(f) $\sigma$ = 0.06
@@
Figure 19: An illustrative example showing different districting plans for Zone 5 in Atlanta. Gray lines represent the basic geographical units patrolled by the police, red lines outline the districting plans. In the map, regions shaded in grey with dashed blue lines represent beats without a dedicated police unit. The presence of blue arrows originating from these beats signifies that their policing needs are addressed by units from adjacent beats.
@@
In Figure 19, we measure the standard deviation of the individual workloads for different overlapping districting plans, denoted as $\sigma$. Panels (a), (b), and (c) display a scenario where a single beat lacks a dedicated police unit. Specifically, in panel (a), the districting plan is non-overlapping since the policing requirements of this unassigned beat are exclusively managed by a single adjacent beat. In panels (b) and (c), we display different overlapping districting plans by varying the number of overlapping beats. In these configurations, the standard deviations of workload are notably reduced, measured at $0.11$ and $0.08$ respectively, demonstrating the impact of overlapping districts on workload distribution uniformity. Panels (d), (e), and (f) depict scenarios wherein two beats operate without dedicated police units. Notably, panel (f) demonstrates a further reduction in workload standard deviation compared to panel (c), measured at $0.06$, attributable to an additional beat being serviced by the overlapping police units from its neighbors.
Plan Beats with Police Unit $\sigma$
@@
In Table 2, we record the workload allocation among police units across beats under different districting plans. Plans (a) – (f) align with corresponding panels in Figure 19. Plan (g) depicts an independent scenario where each beat is serviced by a dedicated police unit with non-overlapping districts. It is noteworthy that the total number of police units varies by districting plan, with plans (a) – (c) utilizing 11 units, plans (d) – (f) utilizing 10 units, and plan (g) utilizing all 12 units. We observe that plan (f) exhibits a lower standard deviation in workload compared to plan (g), despite the latter having two additional police units.
The case study also provides valuable managerial insights for addressing the police redistricting problem. It shows that deploying a larger number of police units does not necessarily equate to more effective policing. For instance, plan (f) with 10 units had a lower workload standard deviation compared to plan (g) with 12 units. This suggests that the utilization of overlapping police beats, as opposed to a rigid, non-overlapping structure, can significantly enhance the efficiency of resource allocation. The overlapping districting approach allows for a more flexible and dynamic response to varying call frequencies, especially in high-demand areas like Zone 5 in Atlanta or other cities that face similar challenges.
@@ 4 Discussions
This paper presents a novel generalized hypercube queue model that captures the queuing dynamics of service systems experiencing heavy traffic. We demonstrate the model’s versatility by applying it to analyze overlapping patrols. Our Markov model employs a hyperlattice to represent an extensive state space, where each vertex corresponds to an integer-valued vector. The transition rate matrix is highly sparse, permitting transitions between vectors with an $\ell_{1}$ distance difference of 1. To enhance computational tractability, we introduce an efficient state aggregation technique that connects our model to a general birth-and-death process with state-dependent rates. This simplifies the detailed balancing equations into a significantly smaller linear system of equations, facilitating the efficient calculation of steady-state distributions. These distributions are essential for evaluating general performance metrics, as acknowledged in Larson’s seminal work.
@@
We emphasize that akin to the rationale in Larson’s original paper [13], no exact analytical expression exists for the hyperlattice model regarding travel time measures for an infinite-line capacity system. This challenge stems from the server’s position not being chosen based on the call’s received region probability distribution during system saturation periods when the server is dispatched consecutively. Consequently, in practice, we must assume that queuing delay is the primary factor influencing the server’s response time for each call.
@@
Additionally, we highlight that our framework concentrates on a random policy controlled by a set of state-dependent matrices, and it can be extended to more intricate decision-making situations. This generalization would empower us to develop and optimize police dispatching policies that take into account queuing dynamics for more realistic systems. This consideration is particularly relevant since conventional First-Come-First-Served (FCFS) approaches may not be optimally efficient, leaving room for enhancement and optimization. Lastly, despite concentrating on police districting in our paper, the generalized hypercube queue model can also be used for modeling mobile servers in a broader range of applications.
Acknowledgement
The authors would like to thank Jonathan Zhou at Georgia Institute of Technology for generating Figures 1, 2, and 3 using Atlanta Police data, and Sarah Heustis-Michell for discussions. The work of Y. Xie is partially supported by an NSF CAREER CCF-1650913, and NSF DMS-2134037, CMMI-2015787, CMMI-2112533, DMS-1938106, and DMS-1830210, and an Atlanta Police Foundation award.
Appendix A Special case: Two servers sharing an overlapping service region
In this section, we discuss a special case of our framework, where the system only allows at most two servers patrolling an overlapping service region, as illustrated by both Figure 5 (a) and (b).
Because only two servers are allows to patrol an overlapping service region, the system can be described by a regular graph, where edges are defined by a set of two vertices. The overlapping service regions can be represented by a pair of server indices ${i,j}$ and the dispatch policy can be simply redefined as $\eta_{ij}$, which represents the probability of choosing server $i$ in overlapping region $(i,j)$ as $\eta_{ij}$ and we have $\eta_{ij}+\eta_{ji}=1$ for $(i,j)\in\mathcal{E}$. Therefore, the upward transition rate defined in (1) can be specified as follows: For an arbitrary state index $u$, the upward transition rate from $B_{u}$ to $B_{u^{i+}}$ is:
$q_{uu^{i+}}=\begin{cases}~{}\lambda_{i}+\sum_{j:n_{j}>0,j\neq i}\eta_{ij}\lambda_{(i,j)},&~{}(a)~{}n_{i}>0,\\ ~{}\lambda_{i}+\sum_{j:n_{j}=0,j\neq i}\eta_{ij}\lambda_{(i,j)}+\sum_{j:n_{j}>0}\lambda_{(i,j)},&~{}(b)~{}n_{i}=0,\end{cases}$ (13)
where (a) represents the rate of assigning a new call to a busy server, while (b) represents the rate of assigning a new call to an idle server.
The following argument justifies the above definition. First, for an arbitrary server $i$, it must respond to a call received in its primary service region $i$. Secondly, we assume a new call will be randomly assigned to a server when both servers are available or there is no available server. Then server $i$ needs to randomly respond to a call in two scenarios: (a) If server $i$ is already busy and the new call is received in the overlapping service region $(i,j)$ when server $j$ is also busy, with the determination made based on the probability $\eta_{ij}$; (b) If server $i$ is currently idle and the new call is received in the overlapping service region $(i,j)$, and server $j$ is either busy or idle (with the determination made based on the probability $\eta_{ij}$).
This redesigned definition assures that Proposition 2 retains its validity. The proof can be found below:
Proof.
To begin with, we define $\bar{\lambda}(B_{u})\coloneqq\sum_{i\in\mathcal{I}}q_{uu^{i+}}$ as the total sum of upward transition rates originating from an arbitrary state $B_{u}$. Given the upward transition rates that leave state $B_{u}$ defined in (1), we have
$\bar{\lambda}(B_{u})=$ $~{}\sum_{i\in\mathcal{I}}q_{uu^{i+}}$
$=$ $~{}\sum_{i:n_{i}>0}q_{uu^{i+}}+\sum_{i:n_{i}=0}q_{uu^{i+}}$
$=$ $~{}\sum_{i:n_{i}>0}\left(\lambda_{i}+\sum_{j:n_{j}>0,j\neq i}\eta_{ij}\lambda_{(i,j)}\right)+\sum_{i:n_{i}=0}\left(\lambda_{i}+\sum_{j:n_{j}=0,j\neq i}\eta_{ij}\lambda_{(i,j)}+\sum_{j:n_{j}>0,j\neq i}\lambda_{(i,j)}\right)$
$=$ $~{}\sum_{i:n_{i}>0}\sum_{j:n_{j}>0,j\neq i}\eta_{ij}\lambda_{(i,j)}+\sum_{i:n_{i}=0}\sum_{j:n_{j}=0,j\neq i}\eta_{ij}\lambda_{(i,j)}+\sum_{i:n_{i}=0}\sum_{j:n_{j}>0,j\neq i}\lambda_{(i,j)}+\sum_{i\in\mathcal{I}}\lambda_{i}$
$=$ $~{}\sum_{i\in\mathcal{I}}\sum_{j\in\mathcal{I}}\mathbbm{1}\{j\neq i\}\left(\mathbbm{1}\{n_{i}>0,n_{j}>0\}\eta_{ij}\lambda_{(i,j)}+\mathbbm{1}\{n_{i}=0,n_{j}=0\}\eta_{ij}\lambda_{(i,j)}\right)$
$+$ $~{}\sum_{i\in\mathcal{I}}\sum_{j\in\mathcal{I}}\mathbbm{1}\{j\neq i\}\mathbbm{1}\{n_{i}=0,n_{j}>0\}\lambda_{(i,j)}+\sum_{i\in\mathcal{I}}\lambda_{i}$
=$\overset{(i)}{=}$ $~{}\sum_{(i,j)\in\mathcal{E}}\left(\mathbbm{1}\{n_{i}>0,n_{j}>0\}\lambda_{(i,j)}+\mathbbm{1}\{n_{i}=0,n_{j}=0\}\lambda_{(i,j)}\right)$
$+$ $~{}\sum_{i\in\mathcal{I}}\sum_{j\in\mathcal{I}}\mathbbm{1}\{j\neq i\}\mathbbm{1}\{n_{i}=0,n_{j}>0\}\lambda_{(i,j)}+\sum_{i\in\mathcal{I}}\lambda_{i}$
=$\overset{(ii)}{=}$ $~{}\sum_{(i,j)\in\mathcal{E}}\left(\mathbbm{1}\{n_{i}>0,n_{j}>0\}\lambda_{(i,j)}+\mathbbm{1}\{n_{i}=0,n_{j}=0\}\lambda_{(i,j)}\right)$
$+$ $~{}\sum_{(i,j)\in\mathcal{E}}\left(\mathbbm{1}\{n_{i}=0,n_{j}>0\}\lambda_{(i,j)}+\mathbbm{1}\{n_{i}>0,n_{j}=0\}\lambda_{(i,j)}\right)+\sum_{i\in\mathcal{I}}\lambda_{i}$
$=$ $\sum_{(i,j)\in\mathcal{E}}\lambda_{(i,j)}+\sum_{i\in\mathcal{I}}\lambda_{i}$
$=$ $~{}\lambda,$
where the equation $(i)$ holds due to $\eta_{ij}+\eta_{ji}=1$ and the equality $(ii)$ also holds because $\lambda_{(i,j)}=\lambda_{(j,i)}$. Hence, the sum of all upward transition rates leaving any given state is constantly $\lambda$.
Now the birth rate of the aggregated state $N_{k},~{}k\geq 0$ can be obtained by multiplying $\bar{\lambda}(B_{u})$ by the number of states that are part of $N_{k}$, which is equivalent to $\binom{k+I-1}{I-1}$. It is also evident that the total of all downward transition rates leading to an arbitrary state is $\sum_{i\in\mathcal{I}}\mu_{i}$, which is equal to $\mu$. Consequently, we can obtain the death rate of $N_{k},~{}k\geq 1$ similarly by multiplying the total service rate $\mu$ by the number of states included in $N_{k}$. ∎

### v4

@@ 4 Experiments
In this section, we evaluate the performance of the hyperlattice model using the proposed metrics in Section 3.4. Our assessment includes a simulated environment in Section 4.1 and a real-world case study on police redistricting in Atlanta, Georgia, conducted in collaboration with the Atlanta Police Department in Section 4.2.
@@ 4.1 Synthetic results
We first evaluate the effectiveness of the proposed hyperlattice queueing model through synthetic experiments. These experiments employ diverse parameter settings within synthetic service systems to investigate the accuracy of the performance metrics estimated by our models. To embark on this, we first use the hyperlattice queueing model to estimate the steady-state probabilities of the systems by solving the balance equations (6). Based on the solutions, we then calculate summary statistics, including individual workloads and the fraction of dispatches, as shown in (10) and (11). To validate the accuracy of our model estimations, we also developed a simulation program using simpy (Prountzos et al. 2021), a process-based discrete-event simulation framework based on Python. The primary purpose of this simulation program is to provide a ground truth for the queueing dynamics of the system, allowing us to compare the estimated performance measures from our model with the simulation results.
The simulation setup involves a $100\times 100$ grid, where the width and height are both set to $100$ units. We explore different levels of overlapping service areas by varying the overlap width from 0 to $99$ in $100$ increments. For each overlap width, we calculate the subregion radius and set the positions of the two servers accordingly. The servers are positioned symmetrically along the $x$-axis at coordinates determined by the radius of the subregion, and each server is responsible for a specific subregion polygon. When an event occurs, its location is randomly determined within the region. The simulation then identifies the police units responsible for the event based on their assigned regions. If multiple police units are available, a dispatch strategy is employed to select which unit will respond. The selected police unit travels to the event location, and the travel time is computed based on the Euclidean distance between the unit’s current position and the event location. Upon arrival, the police unit handles the event, which involves a service time drawn from an exponential distribution based on the unit’s service rate. The police units have a service rate of $1$ and move at a constant speed of 10 units, ensuring that queuing delay remains the predominant factor influencing the server’s response time for each incident. The event arrival rate is set to $1$, with each simulation run lasting $1000$ events. Using the simulation results, we can approximate the true individual workload of each server $\rho_{i}$ by calculating the percentage of time that the server is busy, and the true fraction of dispatches $\rho_{i,e}$ (or $\rho_{i,i}$) by calculating the percentage of calls received in region $e$ (or $i$) that have been assigned to server $i$.
The experiments were conducted on a personal laptop running macOS, equipped with an M3 Pro chip featuring a $12$-core CPU with a clock speed of up to $4.06$ GHz and $18$ GB of LPDDR5 RAM with a memory speed of up to $153.6$ GB/s.
@@ 4.1.1 Overlapping patrolling
We also test our models on the synthetic systems that allow overlapping patrolling. Adapting the truncated hyperlattice approximation from Section 3.3 with an upper bound state of $K=10$, we create a two-server system where the service region is a 100 by 100 square with a single overlapping service region in the middle ($I=2$ and $E=1$), as depicted in Figure 3.1 (a). We assume that calls are uniformly distributed with an arrival rate of $\lambda=1$ and that the service rate of each server is fixed at $\mu_{i}=1$. Changing the hyperparameters of the system, such as the size of the overlapping region and the dispatch policy, can significantly shift the queueing dynamics. Our numerical experiments suggest that the hyperlattice queuing models can accurately capture intricate queuing dynamics under different hyperparameter settings. In Figures EC2 (a) and (b), we observe that the approximation error of both the individual workloads and the fraction of dispatch by the hyperlattice model is less than $0.1$ across all $\lambda/\mu$ ratios. Additionally, in Figure EC2 (c), we show that the mean response time of the system is also closely approximated by the hyperlattice model.
@@
Synthetic results for two-server systems with varying overlapping ratio $r$. In (a), (b), and (c), the shaded area represents the overlapping region of the system. (d) and (e) show individual workloads $\rho_{i}$ and fraction of dispatches $\rho_{i,(i,j)}$ with varying overlapping ratio $r$. The solid lines represent the estimated performance measures using hyperlattice queueing models, whereas the dashed lines represent the true performance measures obtained from simulation runs.
To examine the impact of the overlapping ratio on system dynamics, we first create 100 systems by varying the overlapping ratio $r$. This ratio represents the percentage of the area of the overlapping region with respect to the area of the entire service region of server 1, i.e., $r=|\mathcal{O}_{\{1,2\}}|/(|\mathcal{O}_{\{1,2\}}|+|\mathcal{P}_{1}|)$, as shown in Figure 4.1.1 (a-c). We then compare the estimated values of $\rho_{i}$ and $\rho_{i,e}$ to their true values (indicated by dashed lines) in Figure 4.1.1 (d) and (e), and find that the estimated values are a good approximation of their true counterparts. Clearly, the workloads of the two servers differ more as the overlapping ratio increases, thereby enlarging the area that server 1 has to cover. Moreover, as the ratio increases, the probability of dispatching either server to the overlap region grows (indicated by green and orange lines). In contrast, the probability of sending server 2 to its primary area remains the same (blue line), and the chance of dispatching server 1 decreases (red line). This observation arises from our simplified synthetic setup, where we use a uniform dispatch policy and the dispatch probability to a region directly correlates with its size.
@@
Synthetic results for two-server systems with varying dispatch policy $\eta_{(1,2)}(1)$ and $\eta_{(1,2)}(2)$. In (a), (b), and (c), the shaded area represents the overlapping region of the system. (d) and (e) show individual workloads $\rho_{i}$ and fraction of dispatches $\rho_{i,(i,j)}$ with varying overlapping ratio $r$. The solid lines represent the estimated performance measures using hyperlattice queueing models, whereas the dashed lines represent the true performance measures obtained from simulation runs.
@@
Furthermore, we also investigate the impact of dispatch policy by creating another 100 systems with different values of $\eta_{(1,2)}(1)$ and $\eta_{(1,2)}(2)$ (where $\eta_{(1,2)}(1)=1-\eta_{(1,2)}(2)$), while keeping the service regions the same, as shown in Figure 4.1.1 (a-c). For simplicity, we here omit the subscript state $u$ in the notation of dispatch policy $\eta_{e,u}(i),~\forall i\in e$ as we assume it depends on the state $u$ only only when prioritizing idle servers and remains independent of $u$ when all servers are busy, as discussed in Section 3.2.3. Figure 4.1.1 (d) and (e) show that the estimated $\rho_{i}$ and $\rho_{i,e}$ (indicated by solid lines) can closely approximate their true values suggested by the simulation. As we observe, when the dispatch probability for a call from the overlapping region to either server is 0.5, the workload distribution between the two servers is equal, i.e., $\eta_{1,2}(1)=\eta_{1,2}(2)=0.5$.
@@ 4.2 Case study: Police Redistricting in Atlanta, GA
Police workload in Atlanta, Georgia. (a) presents a comprehensive map delineating the various police zones in Atlanta, Georgia, while (b) specifically focuses on Zone 5. The intensity of color shading directly corresponds to the frequency of police call arrivals, providing a visual representation of call distribution across the zones.
@@
In this section, we present a case study focusing on the police redistricting problem within Atlanta, Georgia. Figure 4.2 illustrates the various police zones of the city, with a particular emphasis on the intricate beats of Zone 5. Ideally, each beat would be served by a dedicated police unit, ensuring a focused response within its boundaries. However, this is often not feasible due to a limited number of available police units, primarily attributed to constraints in officer and resource allocation. In instances where a specific beat lacks a dedicated police unit, units from adjacent beats are required to assist. While such support is permissible among units within the same zone, assistance across zones is explicitly prohibited. The depth of color shading in Figure 4.2 is a direct representation of the frequency of police call arrivals, quantified by the arrival rate $\lambda$. These rates are estimated based on an analysis of historical 911 calls-for-service data, specifically collected during the years 2021 and 2022, as detailed in the study of Zhu et al. (2022). The color indicates that Zone 5 experiences the highest frequency of police calls, averaging approximately 18 calls per hour. As a result, our analysis concentrates on examining potential overlapping districting plans specific to the beats in Zone 5. A key performance metric under consideration is the standard deviation of the individual workloads, as previously discussed in Section 3.4.1.
[…]
@@
An illustrative example showing different districting plans for Zone 5 in Atlanta. Gray lines represent the basic geographical units patrolled by the police, and red lines outline the districting plans. In the map, regions shaded in grey with dashed blue lines represent beats without a dedicated police unit. The presence of blue arrows originating from these beats signifies that their policing needs are addressed by units from adjacent beats.
@@
In Figure 4.2, we measure the standard deviation of the individual workloads for different overlapping districting plans using the truncated hyperlattice model in Section 3.3 with an upper bound state of $K=10$, denoted as $\sigma$. Panels (a), (b), and (c) display a scenario where a single beat lacks a dedicated police unit. Specifically, in panel (a), the districting plan is non-overlapping since the policing requirements of this unassigned beat are exclusively managed by a single adjacent beat. In panels (b) and (c), we display different overlapping districting plans by varying the number of overlapping beats. In these configurations, the standard deviations of workload are notably reduced, measured at $0.11$ and $0.08$ respectively, demonstrating the impact of overlapping districts on workload distribution uniformity. Panels (d), (e), and (f) depict scenarios wherein two beats operate without dedicated police units. Notably, panel (f) demonstrates a further reduction in workload standard deviation compared to panel (c), measured at $0.06$, attributable to an additional beat being serviced by the overlapping police units from its neighbors.
Plan Workload of police units $\sigma$
@@
In Table 2, we record the workload allocation among police units across beats under different districting plans. Plans (a) – (f) align with corresponding panels in Figure 4.2. Plan (g) depicts an independent scenario where each beat is serviced by a dedicated police unit with non-overlapping districts. It is noteworthy that the total number of police units varies by districting plan, with plans (a) – (c) utilizing 11 units, plans (d) – (f) utilizing 10 units, and plan (g) utilizing all 12 units. We observe that plan (f) exhibits a lower standard deviation in workload compared to plan (g), despite the latter having two additional police units.
Table 3: Mean travel times for police units across beats under different districting plans.
Plan Mean travel time of police units (minutes) $\bar{\tau}$}
1 2 3 4 5 6 7 8 9 10 11 12
a 12.15 14.24 - 12.26 12.07 12.20 11.81 11.94 12.01 12.09 12.16 11.82 12.54
b 13.89 13.64 - 12.26 12.07 12.20 11.81 11.94 12.01 12.09 12.16 11.82 12.59
c 13.24 13.10 - 13.75 12.62 12.20 11.81 11.94 12.01 12.09 12.16 11.82 12.62
d 12.15 14.24 - 12.26 12.07 12.20 11.81 11.94 12.01 12.69 - 11.82 12.61
e 13.90 13.64 - 12.26 12.07 12.20 11.81 12.41 12.01 12.44 - 11.82 12.67
f 13.24 13.10 - 13.75 12.62 12.20 12.21 12.19 12.01 12.28 - 11.99 12.63
g 12.15 12.03 12.21 12.26 12.07 12.20 11.81 11.94 12.01 12.09 12.16 11.82 12.07
In Table 3, we present the mean travel times of different police units under various districting plans, as well as the overall mean travel time $\bar{\tau}$}. It is observed that enabling overlapping service regions results in increased mean travel times for police units serving neighboring areas. Although plan (f) exhibits a lower standard deviation in workload compared to plan (g) with fewer police units, it increases the mean travel time for neighboring units by more than one minute and the overall mean travel time by about half a minute.
The case study also provides valuable managerial insights for addressing the police redistricting problem. It shows that deploying a larger number of police units does not necessarily equate to more effective policing. For instance, plan (f) with 10 units had a lower standard deviation of workload compared to plan (g) with 12 units. This suggests that utilizing overlapping police beats, as opposed to a rigid, non-overlapping structure, can significantly enhance the efficiency of resource allocation. However, it should be noted that utilizing overlapping police beats may come at the cost of increased mean travel times. Practitioners must carefully balance the trade-off between response time and resource allocation efficiency in their actual implementations. In general, the overlapping districting approach is preferred in high-demand areas, such as Zone 5 in Atlanta or other cities facing similar challenges.
@@ 5 Discussions
This paper presents a novel generalized hypercube queueing model that captures the detailed queueing dynamics of each server in a system. We demonstrate the model’s versatility by applying it to analyze overlapping patrols. Our Markov model employs a hyperlattice to represent an extensive state space, where each vertex corresponds to an integer-valued vector. The transition rate matrix is highly sparse, permitting transitions between vectors with an $\ell_{1}$ distance difference of 1. To enhance computational tractability, we introduce an efficient state truncation technique that connects our model to a general birth-and-death process with state-dependent rates. This simplifies the detailed balancing equations into a significantly smaller linear system of equations, facilitating the efficient calculation of steady-state distributions. These distributions are essential for evaluating general performance metrics, as acknowledged in Larson’s seminal work.
@@
We emphasize that, akin to the rationale in Larson’s original paper (Larson 1974), no exact analytical expression exists for the hyperlattice model regarding travel time measures for an infinite-line capacity system. This challenge arises from the server’s position not being chosen based on the received region probability distribution of the call during system saturation periods when the server is dispatched consecutively. Consequently, in practice, we must assume that queueing delay is the primary factor influencing the server’s response time for each call.
@@
Additionally, we highlight that our framework concentrates on a random policy controlled by a set of state-dependent matrices, and it can be extended to more intricate decision-making situations. This generalization would empower us to develop and optimize police dispatching policies that take into account queueing dynamics for more realistic systems. This consideration is particularly relevant since conventional First-Come-First-Served (FCFS) approaches may not be optimally efficient, leaving room for enhancement and optimization. Lastly, although our paper focuses on police districting, the generalized hypercube queueing model can also be applied to model mobile servers in a broader range of applications.
An important consideration emphasized in the literature is the explicit modeling of travel times in queueing systems, as region-dependent service rates and spatial dynamics critically influence queue behavior and dispatch effectiveness (e.g., Kanoria (2021), Besbes et al. (2022)). In our hyperlattice queueing model, travel times are treated as exogenous — consistent with Larson’s hypercube formulation — to maintain analytical and computational tractability. Nevertheless, incorporating travel-time dynamics directly into the Markov chain would improve accuracy for spatially sensitive dispatch policies. As an intermediate step, our supplementary experiments in Section EC2 incorporate heterogeneous service rates that reflect server-specific travel times, thereby validating the robustness of the truncated hyperlattice approximation. Extending the model to endogenize travel-time dynamics remains a promising direction for future research.